# Balancing Performance and Efficiency: A Multimodal Large Language Model Pruning Method based Image Text Interaction

### v1

@@ Introduction
From the experimental results, it is evident that under the training-freed setting, our proposed method achieves the best performance. In terms of efficiency, our method averages 127 visual tokens across the six datasets. Under the fine-tuning setting, our method reaches optimal performance on the ScienceQA, POPE, and MMBench datasets, and competitive performance on the VQAv2, TextVQA, and MME datasets. Compared to LLaVA-PruMerge+, our method offers higher computational efficiency (visual tokens: 127 vs 144), and since we also prune text tokens, we hold an advantage in the total token count. In comparison to CrossGET, our method also demonstrates efficiency advantages. Due to fine-tuning enabling the model to adapt to pruned visual tokens, it may achieve better performance on tasks like ScienceQA and MME datasets. Overall, our proposed method strikes a balance between performance and efficiency, showcasing strong competitiveness and potential.
@@
To further evaluate the computational efficiency of our proposed method, we employed open-source simulation software (Yuan et al. 2024) to simulate the computational efficiency on an NVIDIA A100 GPU. During the reverse simulation calculations, we assumed text tokens as input and the average number of visual tokens used by each method across the six datasets as the input visual token count. The experimental results are shown in Table 6.
@@
Conclusion
In this paper, we propose a multimodal large model dynamic pruning method for text-visual interaction. Firstly, we analyze the similarity of CLS visual tokens in the visual encoder, which exhibits a long tail distribution. Therefore, we design an objective function to segment the head and tail in the long tail distribution to extract important visual tokens. Subsequently, the similarity between all tokens is calculated at the LLM layer, and the importance score of each token is obtained by summing up the similarity between the current token and other tokens. Tokens with high importance scores are retained, while tokens with low importance scores are discarded to complete interactive pruning.
The experimental results on multiple datasets showed that our method achieved the best performance on most datasets and maintained a balance between performance and computational efficiency. This provides ideas and methods for future research on efficient multimodal large models
Limitation

### v2

@@ 4.4 Main Result
We conducted comparative experiments with other MM-LLMs and existing MM-LLM token pruning methods. The results are shown in Table 5. The experimental results show that under the training-freed setting, our proposed method performs best with five base models. As the model has not been fine-tuned to adapt to the pruned tokens, in order to balance performance and efficiency, our average compression rate has reached 22% 80% (different datasets and models have varying compression rates). Under the fine-tuning setting, our method reaches optimal performance on the ScienceQA, POPE, and MMBench datasets, and competitive performance on the VQAv2, TextVQA, and MME datasets with the 7.5 times compression ratio. Compared to LLaVA-PruMerge+, our method offers higher computational efficiency (visual tokens: 127 vs 144), and since we also prune text tokens, we hold an advantage in the total token count. In comparison to CrossGET, our method also demonstrates efficiency advantages.
Fine-tuning enables the model to adapt to pruned visual tokens, and it may achieve better performance on tasks like ScienceQA and MME datasets. Overall, our proposed method strikes a balance between performance and efficiency, showcasing strong competitiveness and potential. Compared with Recovery Compression, we avoid the situation of excessive compression by using more tokens to ensure that the performance of the model does not significantly decrease in various tasks.
@@
To further evaluate the computational efficiency of our proposed method, we employed open-source simulation software (Yuan et al., 2024) to simulate the computational efficiency on an NVIDIA A100 GPU. During the reverse simulation calculations, we assumed text tokens as input and the average number of visual tokens used by each method across the six datasets as the input visual token count. The experimental results are shown in Table 6.
@@ 4.5 Visualization
5 Conclusion
In this paper, we propose a dynamic optimization algorithm for multimodal large language models token compression targeting long-tail distribution. Firstly, we analyze the similarity of CLS visual tokens in the visual encoder, which exhibits a long tail distribution. Therefore, we design an objective function to segment the head and tail in the long tail distribution to extract important visual tokens. Subsequently, the similarity between all tokens is calculated at the LLM layer, and the importance score of each token is obtained by summing up the similarity between the current token and other tokens. Tokens with high importance scores are retained, while tokens with low importance scores are discarded to complete interactive pruning.
The experimental results on multiple datasets showed that our method achieved the best performance on most datasets and maintained a balance between performance and computational efficiency. This provides ideas and methods for future research on efficient multimodal large models.
6 Limitation
@@
Appendix A Ablation Study
A.1 Efficiency for Different Compression Ratios
We conducted compression ratios ablation experiments on LLaVA-1.5 and InternVL2, and the experimental results are shown in Table 7 and Table 8. The experimental results indicate that a compression ratio of around 20% is a suitable choice, as it can significantly reduce visual tokens while ensuring that performance does not decrease substantially.
| Compression Ratio | POPE | TextVQA | MME |
| 10% | 78.72 | 54.28 | 1330.63 |
| 20% | 83.35 | 55.13 | 1415.71 |
| 50% | 85.70 | 54.99 | 1486.40 |
| Compression Ratio | POPE | TextVQA | MME |
| 10% | 85.86 | 58.05 | 1345.35 |
| 20% | 86.69 | 59.50 | 1346.96 |
| 50% | 86.87 | 63.84 | 1347.07 |